OpenStudy (anonymous):

factor 8p^3+27q^3 64p^3-1 -23xy+20x^2y^2+6 5c^100-80d^100 (x-p)^2-p^2

OpenStudy (anonymous):

easy, post them in 5 question so i can get 5 medals

OpenStudy (anonymous):

um couldnt u just tell me the answers or just the steps to finish because i have 60 problems like those and im trying to go to bed

OpenStudy (anonymous):

ok mark me as best!

OpenStudy (anonymous):

8p^3+27q^3 =(2p+3q)(4p^2-6pq+9q^2)

OpenStudy (anonymous):

64p^3-1 =(4p-1)(16p^2+4p+1)

OpenStudy (anonymous):

you need the rest

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

want to try the others

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok let me know if you get a problem

OpenStudy (anonymous):

were did u get the -6pq from

OpenStudy (anonymous):

in the first problem

OpenStudy (anonymous):

3*2 8 = 2^3 27 = 3^3